OpenStudy (anonymous):

Medal will be given! A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Btw, I think have have part A completed.

OpenStudy (anonymous):

But I need help with part B, and need to make sure part A is done correctly.

OpenStudy (campbell_st):

multiply the 1st equation by -3 and then add the equations, this will eliminate x and allow you to solve for y

OpenStudy (anonymous):

@campbell_st do you mean divide?

OpenStudy (anonymous):

To eliminate x, shouldn't I preform the opposite operation?

OpenStudy (campbell_st):

well if you multiply the 1st equation by -3 you get and add the 2nd 9x - 21y = 48 -9x + 5y = 16 --------------- -16y = 64

OpenStudy (anonymous):

So that's the correct answer for Part A? I thought it was -3x + 7y = -16 -9x + 5y = 16 -9x + -3x = -12x 7y + 5y = 12y -16 + 16 = 0 -12x + 12y = 0

OpenStudy (anonymous):

But than again, I suck at math.

OpenStudy (anonymous):

And I'm also a bit confused. Where did you get "9x - 21y = 48"?

OpenStudy (campbell_st):

I don't really understand the question.... to me... you are being asked to solve simultaneously in doing so, y = -4 and x = -4 so you have the solution... so perhaps is double the 1st equation and add the 2nd equation and you get -6x + 14y =-32 -9x + 5y = 16 --------------- -15x + 19y = -16 so substituting the solutions x = -4 and y = -4 -15(-4) + 19(-4) = 60 - 76 = -16 so the solution of x = -4 and y = -4 is also a solution to the new equation that's been created.

OpenStudy (anonymous):

So the sum of one of the equations is -4?

OpenStudy (anonymous):

And if so, which one?

OpenStudy (anonymous):

@campbell_st Wait, I think I get it now. Thank you so much! :)